# CMSO-transducing tree-like graph decompositions

### v2

@@ 5.3 Application to bi-join decomposition
Bi-joins were introduced in [18, 25] as a generalisation of modules and splits in undirected graphs and we follow their notation. A bi-join in an undirected graph $G$ is a bipartition $\{X,Y\}$ of $V(G)$ such that $\left|X\right|$, $\left|Y\right|\geq 1$ and there are subsets $X^{\prime}\subseteq X$ and $Y^{\prime}\subseteq Y$ such that $X^{\prime}\times Y^{\prime}\cup(X\setminus X^{\prime})\times(Y\setminus Y^{\prime})\subseteq E(G)$, and there are no additional edges between $X$ and $Y$ in $G$.
@@
As for splits, a bi-join $\{X,Y\}$ is considered trivial if $\left|X\right|=1$ or $\left|Y\right|=1$, and a graph is considered prime wrt bi-join if it has only trivial bi-joins. Let us denote by $\mathcal{B}_{\mathsf{join}}^{G}$ the set of bi-joins of a graph $G$. A proof of the bipartitiveness of bi-joins can be found in [25].
Theorem 5.13 ([25]).
@@
For every undirected graph $G$, the system $\mathcal{B}_{\mathsf{join}}^{G}$ is bipartitive.
A corollary of Theorem 5.13 and Lemma 5.2 we get the existence of a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction taking as input an undirected graph $G$ represented as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$ and outputting the laminar tree representing the set of strong bi-joins of $G$. We would like however, as in the case of splits, to output a relational structure that not only represents the laminar tree induced by $\mathcal{B}_{\mathsf{join}}^{G}$, but also allows to $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduce $\mathbb{G}$. Let’s now explain how to construct such a structure from the laminar tree induced by $\mathcal{B}_{\mathsf{join}}^{G}$. There are two ways to compute such a structure and both are similar to the one for the enhanced canonical split decomposition. Let’s first explain how to compute the one from [17].
Transducing the Skeleton graph [17]
If $X$ is a subset of $V(G)$, let $\equiv_{X}$ be the relation on $X\times X$ where $x\equiv_{X}y$ if $N(x)\cap V(G)\setminus X=N(y)\cap V(G)\setminus X$. One easily checks that $\equiv_{X}$ is an equivalence relation and, if $\{X,Y\}$ is a bi-join, then $\equiv_{X}$ and $\equiv_{Y}$ both have at most two equivalence classes. Let $T$ be the bipartitive tree of $\mathcal{B}_{\mathsf{join}}^{G}$, and for a node $u$, let $u_{1}$ and $u_{2}$ be the, at most two, equivalence classes of $\equiv_{L(T_{u})}$. For every node $u$ of $T$, let’s denote by $G_{u}$ the graph whose vertex set is $\{v_{1},v_{2}\mid v$ is a neighbor of $u\}$, and there is an edge $z_{i}z_{j}$ in $G_{u}$ if there is $x\in z_{i}$ and $y\in z_{j}$ such that $xy\in E(G)$. The skeleton graph associated with $T$ is the graph obtained by taking first the disjoint union of the graphs $G_{u}$, for all nodes $u$ of $T$, and then adding t-edges between copies of same equivalence classes. It is worth mentioning that the skeleton graph is uniquely defined from $T$, and since $T$ is unique, up to isomorphism, one can conclude that the skeleton graph is unique, up to isomorphism. It is routine to check the following.
Lemma 5.14 ([17]).
Let $G$ be a connected graph and let $H$ be its skeleton graph. Then the following hold:
  1. 1.
@@
    two vertices $x$ and $y$ of $G$ are adjacent if and only if there is an alternating path between $x$ and $y$ in $H$;
  2. 2.
    the bipartitive tree of $\mathcal{B}_{\mathsf{join}}^{G}$ can be obtained from $H$ by contracting the graphs $G_{u}$, except vertices of $G$, and considering the $\mathsf{t}$-edges as its set of edges.
We are now ready to transduce the skeleton graph of a graph, the proof of which is similar to the one for the canonical split decomposition.
Theorem 5.15.
@@
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that for any connected graph $G$ represented as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $\tau(\mathbb{G})$ is non-empty and every output in $\tau(\mathbb{G})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ representing the skeleton graph of $G$.
Proof 5.16.
@@
Let $G$ be a connected graph represented by the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $(V(G),\mathcal{B}_{\mathsf{join}}^{G})$ the bipartitive set system consisting of all bi-joins of $G$ represented by the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$, $H$ its skeleton graph represented by the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ and $T$ represented by the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate}\}$-structure, the bipartitive tree of $(V(G),\mathcal{B}_{\mathsf{join}}^{G})$.
@@
Our first step is to transduce $\mathbb{B}$ from $\mathbb{G}$. For this it is sufficient to observe that we can easily define an $\operatorname{MSO}$-formula $\phi_{\mathsf{BIPART}}(X,Y)$ with two monadic free-variables $X$ and $Y$ which defines bi-joins $\{X,Y\}$ of $G$.
@@
Since $\mathbb{B}$ is bipartitive we can apply Lemma 5.2 to transduce $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ from $\mathbb{G}\sqcup\mathbb{B}$. In the following we describe how to transduce $\mathbb{H}$ from $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$.
Contrary to the case of the canonical split decomposition, we cannot deterministically deduce, from $T$, the four markers associated with the bi-join displayed by $T$. However, since the set $\{L(T_{u})\mid u$ is a node of $T\}$ is a laminar set system, one can use Corollary 3.17 to non-deterministically associate with each node $u$ of $T$ a leaf as done in the proof of Theorem 3.1. For each node $u$, let’s denote by $u_{1}$ this representative. Knowing $u_{1}$ and $u$, one can make a copy of $u$ and relate it to the equivalence of $\equiv_{L(T_{u})}$ not containing $u_{1}$. Now that we can create copies, for each node $u$, representing the two equivalence classes of $\equiv_{u}$, we leave, to the reader, the description of the transduction outputting the desired skeleton graph, as one can use the same copying technique as the one in Theorem 5.11.
Transducing a decomposition of its Seidel switch equivalence class [25]
Another way to represent the bipartitive tree and the graph itself is described in [18, 25] which uses graphs characteristic graphs as the ones used to describe a graph with its modular decomposition (see Section 4.2). Again, let $T$ be the bipartitive tree of $\mathcal{B}_{\mathsf{join}}^{G}$. For a node $u$ and $x_{1},\ldots,x_{p}$ leaves in the connected components of $T\setminus u$, not two of them in the same connected component, let $G[x_{1},\ldots,x_{p}]$ be the characteristic graph associated with $u$ and $\{x_{1},\ldots,x_{p}\}$. Unfortunately, while in the case of modular decomposition this graph is unique, and in the case of split decomposition one can choose it in a canonical way in order to describe the edges of $G$, there may be many characteristic graphs that can be associated with $u$. However, they are all degenerate if and only if $u$ is.
Lemma 5.17 ([18, 25]).
Let $G$ be a graph and let $T$ be the partitive tree of $\mathcal{B}_{\mathsf{join}}^{G}$. For every node $u$ of $T$ and every characteristic graph $G_{u}$ of $u$, $G_{u}$ is degenerate if and only if $u$ is degenerate.
One can moreover show that they are equivalent with respect to Seidel switch [25]. For a graph $H$ and $W$ a subset of $V(H)$, the Seidel switch of $H$ at $W$, denoted by $\overline{H}^{W}$, is the graph with $V(H)$ as vertex set and edge set $E(H[W])\cup E(H\setminus W)\cup\{xy\mid x\in W,y\notin W$ and $xy\notin E(H)\}$ [26]. One checks that if $H^{\prime}$ is obtained by a Seidel switch on $W^{\prime}$ in $\overline{H}^{W}$, then $H^{\prime}=\overline{H}^{W\operatorname*{\bigtriangleup}W^{\prime}}$ (see for instance [19, Lemma 11.1]). Now, two graphs $H$ and $H^{\prime}$ are Seidel equivalent if one is isomorphic to a Seidel switch of the other. The following tells that all the graphs in a Seidel equivalece class have the same set of bi-joins.
Lemma 5.18 ([25]).
Let $H$ be a graph and let $W$ be a subset of $V(H)$. Then, $\{X,Y\}$ is a bi-join in $H$ if and only if $\{X,Y\}$ is a bi-join of $\overline{H}^{W}$.
It has been proved that Seidel equivalence classes are in bijection with the set of two-graphs [26]. A two-graph is a pair $(V,\Omega)$ where $\Omega$ is a family of subsets of size $3$ on the finite set $V$ such that each subset of size $4$ of $V$ contains an even number of subsets in $\Omega$. For every undirected graph $H$, one can associate a two-graph $\Omega(H)$ where $\{x,y,z\}$ belongs to $\Omega(H)$ if $x,y$ and $z$ are all pairwise distinct and the number of edges in $H[\{x,y,z\}]$ is odd. Conversely, if $(V,\Omega)$ is a two-graph and $x$ is in $V$, then one can construct a graph $H(\Omega,x)$ with vertex set $V$ and edge set $\{yz\mid\{x,y,z\}\in\Omega\}$. The following is proved in [19, 26].
Theorem 5.19 ([19, Theorem 11.7]).
Let $V$ be a finite set of size at least $3$. Then the following hold:
  1. 1.
    for every two-graph $(V,\Delta)$, the graph $H(\Delta,x)$ is such that $\Omega(H(\Delta,x))=(V,\Delta)$;
  2. 2.
    two graphs $G$ and $H$ with vertex set $V$ are Seidel equivalent if and only if $\Omega(G)=\Omega(H)$.
For two two-graphs $(V_{1},\Omega_{1})$ and $(V_{2},\Omega_{2})$ and $v_{1}\in V_{1}$ and $v_{2}\in V_{2}$, their join wrt $v_{1}$ and $v_{2}$ is the two-graph $(V_{1}\cup V_{2}\setminus\{v_{1},v_{2}\},\Omega)$ where $\Omega=\{\{x,y,z\}\subseteq V_{1}\cup V_{2}\setminus\{v_{1},v_{2}\}\mid\{x,y,z\}\in\Omega_{1}\cup\Omega_{2}$ or $\{x,y,v_{1}\}\in\Omega_{1},\{z,t,v_{2}\}\in\Omega_{2}$ for some $t$, or $\{x,y,v_{2}\}\in\Omega_{2},\{z,t,v_{1}\}\in\Omega_{1}$ for some $t\}$.
A corollary of Theorem 5.19 is that if $\{X,Y\}$ is a bi-join of a graph $H$ with $X\times Y\cap E(H)=X^{\prime}\times Y^{\prime}\cup(X\setminus X^{\prime})\times(Y\setminus Y^{\prime})$, if we let $H[X]$ (resp. $H[Y]$) as defined for split decomposition with the marker $v_{X}^{Y}$ (resp. $v_{Y}^{X}$) adjacent to $X^{\prime}$ (resp. $Y^{\prime}$), $\Omega(H)$ is equal to the join of $\Omega(H[X])$ and of $\Omega(H[Y])$ wrt $v_{X}^{Y}$ and $v_{Y}^{X}$. As observed in [25], making $v_{X}^{Y}$ (resp. $v_{Y}^{X}$) adjacent to $X\setminus X^{\prime}$ (resp. $Y\setminus Y^{\prime}$) gives the same two-graph. As a corollary, if we associate with each node $u$ of the bipartitive tree $T$ of $G$ any characteristic graph and join all of them using t-edges as in the split decomposition, one obtains a structure, called Seidel skeleton, representing the Seidel equivalence class of $G$, which can by definition transduces $G$. We can summarise then the construction of the desired structure with the following lemma.
Lemma 5.20.
Let $T$ be the bipartive tree of $\mathcal{B}_{\mathsf{join}}^{G}$ and let $S(G,T)$ be the Seidel skeleton of $T$. If $S(G,T)$ is considered as a split decomposition and $H$ is the graph it defines, then $G$ is Seidel equivalent to $H$.
A direct corollary is the desired transduction.
Theorem 5.21.
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that for any connected graph $G$ represented as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $\tau(\mathbb{G})$ is non-empty and every output in $\tau(\mathbb{G})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ representing the Seidel skeleton of $G$.
Notice that the skeleton graph allows to transduce deterministically $G$ while the structure described here transduces $G$ non-deterministically by guessing the set $W$ such that Seidel switching it gives $G$.
6 Conclusion

### v4

@@ 5.3. Application to bi-join decomposition
Bi-joins were introduced in [dMR05, Rao06] as a generalization of modules and splits in undirected graphs, and were used for instance in [LdMR07] to decide isomorphism in some graph classes of rank-width at most $2$. We follow definitions and notations from [Rao06]. A bi-join in an undirected graph $G$ is a bipartition $\{X,Y\}$ of $V(G)$ such that $\left|X\right|$, $\left|Y\right|\geq 1$ and there are (possibly empty) subsets $X^{\prime}\subseteq X$ and $Y^{\prime}\subseteq Y$ such that $\big{(}X^{\prime}\times Y^{\prime}\big{)}\cup\big{(}(X\setminus X^{\prime})\times(Y\setminus Y^{\prime})\big{)}\subseteq E(G)$, and there are no additional edges between $X$ and $Y$ in $G$. If $X$ is a module in a graph $G$, then $\{X,V(G)\setminus X\}$ is a bi-join with $X^{\prime}=\emptyset$, and if $V(G)\setminus X$ is complete to $X$, then $Y^{\prime}$ is also equal to the emptyset.
@@
As for splits, a bi-join $\{X,Y\}$ is considered trivial if $\left|X\right|=1$ or $\left|Y\right|=1$, and a graph is considered prime with respect to bi-join if it has only trivial bi-joins. Let us denote by $\mathcal{B}_{\mathsf{join}}^{G}$ the set of bi-joins of a graph $G$. A proof of the bipartitiveness of bi-joins can be found in [Rao06].
{thmC}
@@
[[Rao06]] For every undirected graph $G$, the system $\mathcal{B}_{\mathsf{join}}^{G}$ is bipartitive.
A corollary of Section 5.3 and of 23 is the existence of a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction taking as input an undirected graph $G$ represented as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$ and outputting the bipartitive tree representing the set of bi-joins of $G$. We now propose a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction that takes as input the structure $\mathbb{G}\sqcup\mathbb{T}$ where $\mathbb{T}$ is the structure representing the bipartitive tree of $G$ and outputs a structure that is similar to the canonical split decomposition, from which one can reconstruct the original graph in $\operatorname{MSO}$. A corollary of this $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction is that, for fixed positive $k$, we reduce the existence of a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction for computing rank-decompositions of cut-rank at most $f(k)$, for some function $f$, to the existence of a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction for computing rank-decompositions of cut-rank at most $f(k)$ on prime graphs with respect to bi-join.
Transducing the Skeleton graph [dM03]
Let $G$ be an undirected graph. If $X$ is a subset of $V(G)$, let $\equiv_{X}$ be the binary relation on $X\times X$ where $x\equiv_{X}y$ if $N(x)\cap(V(G)\setminus X)=N(y)\cap(V(G)\setminus X)$. One easily checks that $\equiv_{X}$ is an equivalence relation. Now if $\{X,Y\}$ is a bi-join, then $\equiv_{X}$ and $\equiv_{Y}$ both have at most two equivalence classes. We remind that if $T$ is a tree and $t$ is a node of $T$ and $s$ a neighbour of $t$, then $T_{s}^{t}$ is the connected component of TtT-t containing $s$, and $L(T)$ is the set of leaves of $T$.
Let $T$ be the bipartitive tree of the set $\mathcal{B}_{\mathsf{join}}^{G}$ of the bi-joins of $G$. By construction of $T$, for every node $u$ of $T$ and every neighbour $v$ of $u$, $\{L(T_{v}^{u}),L(T_{u}^{v})\}$ is a bi-join, and then $\equiv_{L(T_{v}^{u})}$ and $\equiv_{L(T_{u}^{v})}$ have both at most two equivalence classes. We denote by $\vec{uv}_{1}$ and $\vec{uv}_{2}$ the two equivalence classes of $\equiv_{L(T_{v}^{u})}$ in what follows (we omit $\vec{uv}_{2}$ if there is only one). Observe that if $v$ is a leaf, then $\equiv_{L(T_{v}^{u})}$, with $u$ its unique neighbour, has a single equivalence class reduced to $\{v\}$ and that we define as $\vec{uv}_{1}$. For every internal node $u$ of $T$, let’s denote by $G_{u}$ the graph whose vertex set is $\{\vec{uv}_{1},\vec{uv}_{2}\mid v$ is a neighbour of u}u\}, and there is an edge $z_{i}z_{j}$ in $G_{u}$ if there is $x\in z_{i}$ and $y\in z_{j}$ such that $xy\in E(G)$. Notice that there are $x\in z_{i}$ and $y$ in $z_{j}$ such that $xy\in E(G)$ if and only if every vertex in $z_{i}$ is adjacent to every vertex in $z_{j}$, and thus the graph $G_{u}$, for every node $u$, is unique up to isomorphism. The Skeleton graph associated with $T$ is a triple $(H,F_{1},F_{2})$ where
  • $H$ is a graph with vertex set $\bigcup\limits_{u\in V(T)\setminus L(T)}V(G_{u})$ and
  • edge set $\bigcup\limits_{u\in V(T)\setminus L(T)}E(G_{u})$,
  • $F_{1}$ is the set of edges $\{\vec{uv}_{i},\vec{vu}_{i}\}$, for $i\in[2]$ for which $uv\in E(G)$.
  • $F_{2}$ is the set of pairs $\{\vec{uv}_{1},\vec{uv}_{2}\}$ for every edge $uv$ of $E(T)$.
We call the edges of $H$ $\mathsf{c}$-edges, edges in $F_{1}$ $\mathsf{t}$-edges and edges in $F_{2}$ $\mathsf{r}$-edges and we naturally model the skeleton graph as $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge},\mathsf{r}\text{-}\mathsf{edge}\}$-structures.
It is worth mentioning that the Skeleton graph is uniquely defined from $T$, and since $T$ is unique, up to isomorphism, one can conclude that the Skeleton graph is unique, up to isomorphism. The Skeleton graph defined in [dM03] does not include $\mathsf{r}$-edges, but we need such edges to be able to recover the graphs $G_{u}$ from the Skeleton graph because $G_{u}$ without the $\mathsf{r}$-edges is not necessarily connected. See Figure 9 for an example.
It is routine to show the following.
{lemC}
Let $G$ be a connected graph and let $(H,F_{1},F_{2})$ be its Skeleton graph. Then the following hold:
  1. (1)
@@
    two vertices $x$ and $y$ of $G$ are adjacent if and only if there is a path between $x$ and $y$ in $(H,F_{1},F_{2})$ on which $\mathsf{c}$-edges and $\mathsf{t}$-edges alternate;
  2. (2)
    the vertices of $G$ are precisely the vertices of $(H,F_{1},F_{2})$ that are not incident to any $\mathsf{t}$-edges.
Figure 9. A graph $G$ with three non-trivial strong bi-joins (left), the partitive tree of $\mathcal{B}_{\mathsf{join}}^{G}$ (middle) and its Skeleton graph (right). Dashed edges are $\mathsf{t}$-edges, squiggly edges are $\mathsf{r}$-edges and the rest are $\mathsf{c}$-edges. Equivalence class vertices are represented by squares apart from the singleton equivalence classes corresponding to vertices of $G$. Note that the edges of the graph $G$ correspond one-to-one to paths in the skeleton graph that alternate between $\mathsf{t}$-edges and $\mathsf{c}$-edges. The graph depicted was taken from [Rao06, Figure 4.5].
It is routine to write an $\operatorname{MSO}$-formula checking that, between two vertices, there is a path that alternate $\mathsf{c}$-edges and $\mathsf{t}$-edges. Hence, we can conclude that there is an $\operatorname{MSO}$-transduction which on input of the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge},\mathsf{r}\text{-}\mathsf{edge}\}$-structures $\mathcal{H}$ modeling the skeleton graph of $G$ outputs the $\{\mathsf{edge}\}$-structure $\mathbb{G}$ modeling $G$.
We are now ready to construct the $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction that computes the Skeleton graph of a graph.
Theorem 26.
@@
There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that for any connected graph $G$ represented as the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $\tau(\mathbb{G})$ is non-empty and every output in $\tau(\mathbb{G})$ is equal to some $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge},\mathsf{r}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ representing the Skeleton graph of $G$.
Proof 5.5.
@@
Let $G$ be a connected graph represented by the $\{\mathsf{edge}\}$-structure $\mathbb{G}$, $(V(G),\mathcal{B}_{\mathsf{join}}^{G})$ the bipartitive set system consisting of all bi-joins of $G$ represented by the $\{\mathsf{BIPART}\}$-structure $\mathbb{B}$, $(H,F_{1},F_{2})$ its Skeleton graph represented by the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{c}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ and $T$ represented by the $\{\mathsf{t}\text{-}\mathsf{edge},\mathsf{degenerate}\}$-structure, the bipartitive tree of $(V(G),\mathcal{B}_{\mathsf{join}}^{G})$.
@@
Our first step is to transduce $\mathbb{B}$ from $\mathbb{G}$. For this it is sufficient to observe that we can easily define an $\operatorname{MSO}$-formula $\phi_{\mathsf{BIPART}}(X)$ with one monadic free-variable $X$ which defines bi-joins $\{X,U\setminus X\}$ of $G$.
@@
Since $\mathbb{B}$ is bipartitive, we can apply 21 to transduce $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$ from $\mathbb{G}\sqcup\mathbb{B}$, where $\mathbb{T}$ is the structure representing the bipartitive tree of $G$. In the following we describe how to transduce $\mathbb{H}$ from $\mathbb{G}\sqcup\mathbb{B}\sqcup\mathbb{T}$.
  • The first transduction chooses non-deterministically a node $r$ of $T$ and root $T$ at $r$ and defines the laminar set system $\mathbb{S}$ where $\mathsf{SET}(X)$ holds if there is a node $u$ with $X$ the set of leaves of the subtree of $(T,r)$ rooted at $u$. Let’s denote by $L_{u}$ the set of leaves of the subtree rooted at $u$.
  • By 12 and 13 there are four bi-colourings $(A_{i},B_{i})_{i\in[4]}$ identifying $V(T)$ and for each $i\in[4]$, there are $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formulas ${\mathsf{repr}}_{B_{i}}(a,X)$ and ${\mathsf{repr}}_{A_{i}}(b,X)$ that are satisfied exactly when $X$ is, respectively, $A_{i}$-represented by $a$ and $B_{i}$-represented by $b$. The second transduction guesses non-deterministically the four bi-colouring $(A_{i},B_{i})_{i\in[4]}$ that identifies nodes of $(T,r)$ as a laminar tree of the laminar set system $\mathbb{S}$. Notice we can filter by accepting only the guesses with unique request and such that, for each node, there is a exactly one $i\in[4]$ such that it is uniquely $A_{i}$-represented and uniquely $B_{i}$-represented.
  • The third transduction makes four copies of each internal node $u$, adds $\mathsf{r}$-edges between copies $1$ and $2$, and between copies $3$ and $4$.
  • It remains now to create the $G_{u}$ graphs, for each internal node $u$. Assume $v$ is the parent of $u$, and let $a_{v}$ be the leaf such that (${\mathsf{repr}}_{A_{i}}(a_{v},L_{v})$ or ${\mathsf{repr}}_{B_{i}}(a_{v},L_{v})$) and $a_{v}\notin L_{u}$, and let $a_{u}$ be a leaf such that ${\mathsf{repr}}_{A_{j}}(a_{u},L_{u})$. Notice that $a_{v}$ exists because $v$ is the least common ancestor of the two leaves that $A_{i}$ and $B_{i}$-represent $v$, and if both do not belong to $L_{u}$, we deterministically define $a_{v}$ as the one that $A_{i}$-represents $v$. Copy $1$ of $u$ will play the role of the equivalence class of $\equiv_{T_{v}^{u}}$ containing $a_{v}$. The copy $2$ will play the role of the equivalence class of $\equiv_{T_{v}^{u}}$ not containing $a_{v}$. The copy $3$ will play the role of the equivalence class of $\equiv_{L_{u}}$ containg $a_{u}$, and the copy $4$ will play the role of the equivalence class of $\equiv_{L_{u}}$ not containing $a_{u}$. The vertex set of $G_{u}$ will be copies $1$ and $2$ of $u$, and the copies $3$ and $4$ of each child of $u$. Notice that there is an edge between two equivalence classes $z$ and $z^{\prime}$ when there are vertices $x\in z$ and $y\in z^{\prime}$ such that $xy$ is an edge of $\mathbb{G}$. Since we have access to a vertex representing an equivalence class in $L_{w}$, for $w$ a child of $u$, and a vertex representing an equivalence class in $L(T_{v}^{u})$, one can write a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formula checking that the other equivalence class of $L_{w}$, $w$ a child of $u$, (resp. of $L(T_{v}^{u})$) is non-empty, and then whether there are edges between two equivalence classes. We can therefore write a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formula stating that two copies of vertices are related by a $\mathsf{c}$-edge.
  • The fourth transduction adds the $\mathsf{t}$-edges. We add a $\mathsf{t}$-edge between a copy $i\in\{1,2\}$ of $u$ and a copy $j\in\{3,4\}$ of $u$, if a vertex of the equivalence class the $i$th copy represents is adjacent to a vertex of the equivalence class the $j$th copy represents. Such edges can be $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-defined in the same way the $\mathsf{c}$-edges are defined, except they relate only particular copies of a same node.
The composition of the four transductions above computes the Skeleton graph of $G$. Since the non-determinism comes only from the construction of $\mathbb{T}$ and the guessing of the bi-colouring, and we can filter after each such step, we guarantee that an output always satisfies the desired outcome.
6. Conclusion